Straight Line Method of Amortization Help

Alpha Company purchased a $1,000, 5 years, 6% bond on July 1, 2019 for $940. Interest is paid semi-annually on June 30 and Dec ember 31. The straight line method of amortization is used for both premiums & discounts. Use this information to prepare the adjusting General Journal entry (without explanation) for the six-months ended December 31, 2019. If no entry is required then write "No Entry Required."
Answer #1

Par value of bonds $=\$ 1,000$

Purchase price of bonds $=\$ 940$

Discount on bonds = Par value of bonds - Purchase price of bonds

$=1,000-940$

$=\$ 60$

Semi annual interest receipts on bonds = Par value of bonds $\times$ Interest rate $\times 6 / 12$

$=1,000 \times 6 \% \times 6 / 12$

$=\$ 30$

Semi annual amortization of bond discount = Discount on bonds/ Semi annual interest periods

$=60 / 10$

$=\$ 6$

Date Accounts Debit Credit
12/31/19 Cash $30
Discount on bond investment $6
Interest revenue $36



answered by: anglebay